# On the Robustness of Fully-Spiking Neural Networks in Open-World Scenarios using Forward-Only Learning Algorithms

### v1

@@ Abstract
In the last decade, Artificial Intelligence (AI) models have rapidly integrated into production pipelines propelled by their excellent modeling performance. However, the development of these models has not been matched by advancements in algorithms ensuring their safety, failing to guarantee robust behavior against Out-of-Distribution (OoD) inputs outside their learning domain. Furthermore, there is a growing concern with the sustainability of AI models and their required energy consumption in both training and inference phases. To mitigate these issues, this work explores the use of the Forward-Forward Algorithm (FFA), a biologically plausible alternative to the Backpropagation algorithm, adapted to the spiking domain to enhance the overall energy efficiency of the model. By capitalizing on the highly expressive topology emerging from the latent space of models trained with FFA, we develop a novel FF-SCP algorithm for OoD Detection. Our approach measures the likelihood of a sample belonging to the in-distribution (ID) data by using the distance from the latent representation of samples to class-representative manifolds. Additionally, to provide deeper insights into our OoD pipeline, we propose a gradient-free attribution technique that highlights the features of a sample pushing it away from the distribution of any class. Multiple experiments using our spiking FFA adaptation demonstrate that the achieved accuracy levels are comparable to those seen in analog networks trained via back-propagation. Furthermore, OoD detection experiments on multiple datasets (e.g., Omniglot and Not-MNIST) prove that FF-SCP outperforms avant-garde OoD detectors within the spiking domain in terms of several metrics used in this area, including AUROC, AUPR, and FPR-95. We also present a qualitative analysis of our explainability technique, exposing the precision by which the method detects OoD features, such as embedded artifacts or missing regions, in multiple instances of MNIST and KMNIST. Our results underscore the enhanced robustness that can be achieved by analyzing the latent spaces produced by already trained models.
Keywords Forward-Forward Algorithm  $\cdot$ Forward-Only Learning  $\cdot$ Spiking Neural Networks  $\cdot$ Out-of-Distribution Detection  $\cdot$ Explainable Artificial Intelligence
@@ 1 Introduction
The fast development of learning algorithms has drastically impacted our dependency on Artificial Intelligence (AI) models, as their exceptional capabilities have led to their continual integration into production-ready software. However, this rapid pace has not been accompanied by additional safety guarantees to address well-known issues faced by modern AI systems. Among these shortcomings, the lack of robustness against inputs outside the training distribution poses a significant challenge in ensuring functionality in scenarios where generalization capabilities are essential [1]. Consequently, many people are hesitant to rely on AI for critical decisions in fields such as medicine or law, even if the model’s performance surpasses that of human experts on specific tasks [2]. This limitation on the generalization capability arises from the predominant use of gradient Back-Propagation (BP) for training models within the AI field. Although these models demonstrate excellent performance across a wide range of tasks, their tendency to overfit on training data and their inadequate responses to novel inputs make them highly unreliable tools for achieving high robustness standards.
Motivated by the learning capabilities of the brain, a novel biologically plausible learning heuristic denoted as the Forward-Forward Algorithm (FFA) was recently introduced by Hinton [3], which has showcased competitive accuracy when compared with BP. However, its most promising property lies within its latent space, which provides fundamental characteristics required for developing robust models. As multiple authors have observed since then [4, 5], the emerging latent space of this algorithm is characterized by classes residing tightly packed together in small clusters, while remaining distinct from other classes. Having tight compression of the data is pivotal in ensuring that models behave correctly when presented with data outside the training distribution, as the latent vector representation would lie outside any of the clusters identified from the training latent space. This property is crucial for ensuring a correct behavior with data from training datasets, as samples outside these clusters can be quickly identified as not belonging to the training distribution. Although the algorithm was initially presented as a biologically plausible alternative, most implementations have relied on analog neural models, leading to reduced research on their spiking counterparts. However, the set of papers working on spiking adaptations have shown how FFA can achieve competitive accuracy levels [6], paving the way for further neuromorphic implementations that promise faster and more energy efficient alternatives.
This work presents a seamless integration of FFA into Spiking Neural Networks (SNNs) via surrogate gradients to fully leverage the internal properties of spiking models. By capitalizing on their characteristic latent space, we create a novel Out-of-Distribution (OoD) Detection Algorithm that uses the latent distance between a sample and the set of In-Distribution (ID) clusters to determine whether a sample should be labeled as either OoD or ID. To extend the functionality of this algorithm, we have developed an interpretability technique that can generate an attribution map for a specific input instance with respect to any arbitrary class, highlighting features that do not correspond with those expected from the distribution of the desired class. This algorithm eases the task of detecting possible artifacts that could be affecting the model’s decision among samples labeled as OoD. Our results showcase the enhanced OoD detection accuracy in spiking networks when employing FFA, demonstrating that its enhanced representational properties improve the robustness of the models. Furthermore, our attribution method proves effective in detecting key features that push a data sample from its respective class towards OoD regions.
@@
The rest of this paper is structured as follows: Section 2 provides a comprehensive review of related work. Section 3 elaborates on all the methodological approaches developed in this work. Specifically, Subsection 3.1 presents the approach taken to adapt FFA for SNNs, Subsection 3.2 describes the FF-SCP algorithm, and Subsection 3.3 introduces our attribution algorithm. Next, Section 4 presents the research questions that this work addresses, along with the experimental setup used for the experiments. Subsequently, Section 5 presents the experimental results, along with a discussion on the observed limitations. Finally, Section 6 outlines the conclusions of the paper and delineates directions for future research.
@@ 2 Related Work and Contributions
In this section, we present a comprehensive overview of relevant literature for this work, focusing on the following central topics: Spiking Neural Networks (Subsection 2.1); Forward-only Learning Algorithms (Subsection 2.2); Out-of-Distribution Detection (Subsection 2.3); and Explainability and Attribution (Subsection 2.4). Finally, Subsection 2.5 details the contribution that this papers offer with respect to the current literature.
@@ 2.1 Spiking Neural Networks
SNNs constitute a special category of artificial neural networks inspired by biological models, aiming to replicate the communication patterns found in real neurons within the brain. Unlike traditional approaches, primarily driven by Analog Neural Networks (ANNs) that rely on continuous activation values to forward information, spiking models operate by using discrete pulses of information, known as spikes. The internal dynamics of the neuron, determined by the chosen neural model, dictate when spikes occur and are typically presented as a dynamical system that provides a mathematical approximation of the behavior observed in real neurons [7]. Consequently, neurons can exhibit highly complex dynamics depending on the neural model, the degree of biological plausibility and the total computational complexity desired for the simulation.
In this work, we rely on the Leaky Integrate-and-Fire (LIF) neural model, known for preserving the standard action-potential dynamic while achieving high computational efficiency [8, 9]. Under this model, the membrane potential $U$ of the neuron is tracked and updated in response to the external stimuli $I_{\text{in}}$, which represents the information forwarded to the neuron via the axon. Whenever the membrane potential reaches a threshold $\theta_{\text{thr}}$, a spike is generated and forwarded to subsequent neurons, at which point the membrane potential is reset to zero. To prevent unbounded activity, a constant decay $\tau$ is applied, exponentially reducing the membrane potential over time. Under a discrete time formulation, the neuron’s membrane potential $U(t)$ is updated as follows:
$U(t)=\tau\cdot U(t-1)+R\cdot I_{\text{in}}(t),$ (1)
where $U(t-1)$ represents the membrane potential at the previous time step, and $R$ is a scaling factor of the input. Once $U(t)$ is computed, the value of the output spike $S(t)$ is determined by the activity of the membrane:
@@
By relying on these models, the neuromorphic computing field has been a highly active area due to the energy and speed advantages they offer, demonstrating proficiency in multiple artificial intelligence tasks. In the field of Computer Vision, deep spiking neural networks have been effectively employed for image classification [10] and image detection tasks [11]. In the Natural Language Processing domain, Zhu et al. implemented a fully spiking version of a Large Language Model, demonstrating its competitiveness against analog language models such as GPT-2 [12]. Additionally, notable advantages have been observed in a brain-like navigation systems by utilizing LIF neurons trained by Spike-Timing-Dependent Plasticity (STDP), a biologically plausible learning rule [13].
@@ 2.2 Forward-only Learning Algorithms
BP stands as the most commonly employed algorithm for training artificial neural networks. This method computes exact gradient values of the weights with respect to a loss function by back-propagating the error from the output through the layers using the chain rule. Due to the precision of weight updates, BP has achieved state-of-the-art results in terms of accuracy across a multitude of tasks. However, the requirement for differentiability in the functions defining the network poses a clear limitation for non-differentiable neural models, such as those in spiking systems. To address this limitation, the most commonly employed approach involves replacing the neural model with a differentiable approximation so that the backward pass can create a faithful approximation of the gradients [14]. Additionally, due to the optimization nature of the problem, models trained with BP have been observed to be prone to overfitting [15], consequently failing to generalize beyond the training distribution and resulting in unknown outcomes when presented with new, unseen data.
To overcome the present limitations of BP, the NeuroAI field has recently emerged as a strong alternative movement. This field advocates for incorporating and studying learning mechanisms observed in the brain to improve current models, aiming for fast learning and highly robust learning models [16]. Among these approaches, the branch of forward-only algorithms poses a viable method for bridging the gap between neuroscience and AI by focusing on a reduced set of biological properties while preserving the advantage of being computationally efficient and staying close to traditional AI methods. These algorithms specifically address the weight transport and update lock problems by relying solely on forward passes. By avoiding these biological incompatibilities, these algorithms also enhance integrability into neuromorphic systems, which are better suited for spiking models.
@@
Recently, the Forward-Forward Algorithm was introduced as a novel forward-only alternative to BP, advocating for a biologically motivated heuristic in which the backward pass is replaced by an additional contrastive forward pass [3]. In this secondary pass, the model is presented with an adversarial sample to contrastively learn to distinguish between samples from the training set (referred to as positive or $D_{\oplus}$) and synthetically created data (referred to as negative or $D_{\ominus}$). As it follows from to its forward-only nature, each layer is trained independently, using a layer-specific loss function that relies solely on local information. To formulate this loss function, Hinton proposed the use of a goodness score, which measures the likelihood of a latent vector from belonging to the positive data distribution. This goodness function $G:\mathbb{L}\rightarrow\mathbb{R}$ is defined as the squared Euclidean distance of the latent vector from its original position, where $\mathbb{L}$ denotes the latent space of the given layer. This choice was motivated by the simplicity of its derivative, which facilitates clear and effective weight update expressions. Under this formulation, the goodness score serves as a measure of the layer’s activity, advocating for achieving highly active layers when presented with positive samples, and nearly dormant outputs when encountering a negative instance.
To avoid relying on a regression-based task, which could provoke unstable learning dynamics, a probability-based approach was taken. By using the Binary Cross-Entropy loss with a probability function defined over the goodness scores, the approach ensures a balanced distribution of positive and negative activities. The probability function $P:\mathbb{R}\rightarrow[0,1]$ is chosen to map the goodness scores to a final probability estimation, indicating the likelihood of a latent vector being obtained from a sample from the positive data distribution. This function is defined by the logistic function, modulated by a threshold value $\theta$ to shift the scores toward zero when presented with near-zero goodness scores. The formal expression for this function is given by $P(x;\theta)=\text{sigmoid}(x-\theta)$.
@@
As previously stated, each learning step involves two contrastive forward passes: a positive pass, where each layer is presented with a sample from the training distribution and optimized to maximize its goodness score; and a negative pass, where a synthetic sample is employed to teach the model to remain dormant when presented with non-training data. In its supervised formulation, positive instances are created by embedding a one-hot encoding vector of the corresponding label into the corner of the base input image. Analogously, negative data is created by taking a sample from the training dataset and embedding a random label that does not align with its corresponding one. Using this positive and negative generation method, the inference process is achieved by computing the goodness score of the data sample with respect to all labels, thereby selecting the label achieving the highest score as the output class. In a subsequent work, Lee et al. proposed a refinement of this method by using random binary vectors associated with each class, which were then appended to the end of the vector of the data sample, reducing the overlap between the data and the label, and resulting in improved classification accuracy [17].
The algorithm has gained significant attention from NeuroAI researchers, leading to multiple improvements over the original formulations, which have resulted in increased accuracy [18, 17] and improved biological plausibility [5, 6]. Expanding on the latter topic, Ororbia and Mali pioneered additional biological heuristics for FFA by integrating predictive coding, which allows the model to provide accurate reconstructions of the latent space [5]. Simultaneously, Ororbia proposed the first implementation of an event-driven version of FFA, replacing the conventional error-driven update with modulated Hebbian learning rules, demonstrating a straightforward adaptation of FFA into a biologically plausible spiking learning pipeline [6]. Both of these works, along with [4] and [19], have presented consistent results showcasing the representational properties that emerge from this algorithm. Their work highlights the particular latent topology that results after training, characterized by sparse latent activation and neural specialization that leads to several compact clusters comprised of instances of the same class, which lie far from each other, rendering class clusters highly separated within the space.
@@ 2.3 Out-of-Distribution Detection
OoD detection is a crucial task for developing production-ready software, as it ensures that models are not presented with inputs outside their training distribution, which could result in unpredictable and potentially dangerous outputs [1]. To achieve this, OoD detection algorithms define a scoring function that measures the likelihood of a sample from belonging to the ID dataset, which is then used to label input data as either ID or OoD depending on the score obtained. When studying how OoD samples emerge, it’s important to distinguish between two fundamentally different scenarios: semantic shifts, which occur when samples that do not belong to any class start to emerge, changing the overall distribution of the input space without truly altering the distribution of the previously seen classes; and covariate shifts, which occur when the inherent distribution of one or several classes shifts from its original distribution. In this context, OoD detection only covers the detection of semantic shifts in the data, as covariate shifts are covered by the branch of anomaly detection [20]. Based on the review from Yang et al. [1], OoD detection algorithms can be categorized into four distinct heuristics based on the information used to detect OoD samples: classification-based, distance-based, density-based, and reconstruction-based methods. Our paper focuses on distance-based techniques, which rely on the distances within the latent space to define the desired OoD scoring function. Due to the lack of literature on OoD detection in networks trained using the FFA, we will provide a comparative perspective on methods employing backpropagation BP.
Currently, the ODIN algorithm stands as the most widely-used benchmark for OoD detection due to its high accuracy and reduced computational costs. This algorithm operates by measuring the effect on the softmax distribution of the last layer when changing the temperature scaling and introducing small perturbations into the inputs, where OoD samples result in the greatest differences [21]. An alternative approach was proposed by Bergman et al., where the deep k-nearest neighbor algorithm is employed to measure the distance between input samples and previously computed ID representative latent manifolds, thereby using the latent distance to detect anomalies [22]. Building upon this research, Sun et al. extended the application of the deep k-nearest-neighbor distance for OoD detection [23], highlighting the importance of exploiting the geometric properties in the latent space of trained networks for OoD detection tasks. Despite the extensive literature on this task, a noticeable gap exists in the context of techniques designed specifically for spiking neural networks. In their recent work, Martinez-Seras et al. introduced a novel method for OoD detection that focuses on spike count patterns in the latent space. By characterizing spike patterns across various layers, they identify whether a given spike pattern aligns with those extracted from in-distribution data [24].
2.4 Explainability and Interpretability
Explainable AI, commonly referred to as XAI, is a field focused on developing techniques to offer humanly understandable explanations for the motivations behind a model’s outputs. It is well acknowledged that most modern models operate as black boxes, achieving high performance across multiple tasks but offering little to no insight into the heuristics they follow to obtain their outputs. To address this shortcoming, the XAI field encompasses multiple subtopics aimed at extracting different types of knowledge of models. Among these subtopics, the most relevant are given by: understandability, comprehensibility, interpretability, explainability, and transparency. This paper focuses on Interpretability, which aims to extract meaningful explanations that can be easily understood by humans agents.
As of now, the most prevalent interpretability techniques for computer vision tasks are driven by Class Activation Maps (CAMs) [25], which create visual insights highlighting the most significant regions of an input image that impacted the prediction. Among methods providing this type of explanation, Grad-CAM stands as the one with the greatest impact. This method uses the gradient information of the last convolutional layer to detect the areas of the input that had the highest impact on the prediction of a specific class [26]. Although several extensions of the algorithm have been developed to enhance the quality of the attribution maps [27, 28], these methods are not applicable to spiking networks due to their reliance on precise gradient information, which cannot be obtained from spiking networks because of the non-differentiability of their neural model. To overcome this issue, Kim et al. introduced a visualization algorithm named Spike Activity Mapping, which leverages the temporal dynamics of SNNs to compute neuronal contribution scores during forward propagation based on the history of spikes [29]. Another approach proposed by Bitar et al. addresses the gradient issue in image attribution by employing surrogate gradients, demonstrating their capability to maintain sufficient accuracy for effectively modeling the saliency map [30].

### v2

# Forward-Forward Learning achieves Highly Selective Latent Representations for Out-of-Distribution Detection in Fully Spiking Neural Networks

@@ Abstract
In recent years, Artificial Intelligence (AI) models have achieved remarkable success across various domains, yet challenges persist in two critical areas: ensuring robustness against uncertain inputs and drastically increasing model efficiency during training and inference. Spiking Neural Networks (SNNs), inspired by biological systems, offer a promising avenue for overcoming these limitations. By operating in an event-driven manner, SNNs achieve low energy consumption and can naturally implement biological methods known for their high noise tolerance. In this work, we explore the potential of the spiking Forward-Forward Algorithm (FFA) to address these challenges, leveraging its representational properties for both Out-of-Distribution (OoD) detection and interpretability. To achieve this, we exploit the sparse and highly specialized neural latent space of FF networks to estimate the likelihood of a sample belonging to the training distribution. Additionally, we propose a novel, gradient-free attribution method to detect features that drive a sample away from class distributions, addressing the challenges posed by the lack of gradients in most visual interpretability methods for spiking models. We evaluate our OoD detection algorithm on well-known image datasets (e.g., Omniglot, Not-MNIST, CIFAR10), outperforming previous methods proposed in the recent literature for OoD detection in spiking networks. Furthermore, our attribution method precisely identifies salient OoD features, such as artifacts or missing regions, hence providing a visual explanatory interface for the user to understand why unknown inputs are identified as such by the proposed method.
keywords:
Spiking Neural Networks, Out-of-Distribution Detection , Forward-Forward Algorithm , Explainable Artificial Intelligence
journal: Neural Networks
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
Surrogate gradients allow for spiking adaptations of the Forward-Forward Algorithm
The topology of Forward-Forward networks splits in and out of distribution samples
Highly selective neurons can produce highly interpretable attribution maps
@@ 1 Introduction
The fast development of learning algorithms has drastically impacted our dependency on Artificial Intelligence (AI) models, whose exceptional capabilities have led to their continual integration into production-ready software. However, this recent surge has been heavily hindered by two main factors: first, the vast computational power required to train these models; and second, the lack of robustness guarantees to ensure that these models function reliably on novel or unexpected data. In contrast to these systems, biological brains operate using a negligible amount of energy, while demonstrating a strong capacity to learn and adapt to different data streams (Zador et al., 2022). Recent research efforts have focused on developing highly parallel event-driven systems, inspired by the spike-driven dynamics of biological neurons, with the goal of providing an energy-efficient alternative to traditional GPUs (Tavanaei et al., 2019). Due to their architectural design, these computers, denoted as neuromorphic computers, operate as a natural physical implementation of Spiking Neural Networks (SNNs), thereby obtaining higher inference speeds. In addition, due to their natural resemblance to biological brains, SNNs facilitate the use and development of biologically-inspired algorithms, which are known to exhibit a natural resistance to Out-of-Distribution (OoD) samples, providing an alternative mechanism to prevent undesired inputs and behavior in networks. One recent algorithm following these biological motivations is the so-called Forward-Forward Algorithm (FFA), which employs a local contrastive-based learning rule to make networks highly selective to In-Distribution (ID) data, while causing OoD samples to produce almost dormant latent activity (Hinton, 2022). This learning heuristic, beyond demonstrating competitive accuracy with respect to backpropagation in dense networks, has also been shown to reproduce the highly representative properties of the latent space of biological models, which segregates the data into feature-selective clusters (Tosato et al., 2023; Ororbia and Mali, 2023). However, although some papers have proposed alternative implementations of this algorithm for SNNs (Ororbia, 2023; Terres-Escudero et al., 2024), the vast majority of research is still focused on the non-spiking original implementation. Similarly, the field of OoD detection is predominantly centered around algorithms developed for analog networks, which are not able to obtain the same quantitative results when adapted for SNNs (Martinez-Seras et al., 2023). In this context, the biological foundation of FFA, which makes it inherently suitable for SNNs, combined with its contrastive learning process that encourages OoD samples to diverge from ID inputs, presents a compelling synergy.
This work relies on a spiking implementation of the FFA to develop a fully-spiking OoD detection algorithm that exploits the geometrical properties of the latent space to define a distance-based scoring function. By providing empirical evidence of the large separation between ID and OoD clusters in the latent space, we propose a heuristic to determine whether an input belongs to the ID latent set. Given the importance of explainability mechanisms in the development of AI based systems (Díaz-Rodríguez et al., 2023), we extend our results by proposing an algorithm that generates an attribution map for a given input, highlighting features that do not align with those expected from the ID samples of any desired class. The results presented in this paper demonstrate that spiking neural networks can achieve competitive robustness against unknown stimuli when integrated with biologically plausible frameworks, positioning them as a energy-efficient and safe modeling approach for responsible and trustworthy AI systems.
@@
The rest of this paper is structured as follows: Section 2 provides a comprehensive review of related work. Section 3 elaborates on all the methodological approaches developed in this work. Specifically, Section 3.1 presents the approach taken to adapt FFA for SNNs, Section 3.2 describes the FF-OoD algorithm, and Section 3.3 introduces our attribution algorithm. Next, Section 4 presents the research questions that this work addresses, along with the experimental setup used for the experiments. Subsequently, Section 5 presents the experimental results, along with a discussion on the observed limitations. Finally, Section 6 outlines the conclusions of the paper and delineates directions for future research.
@@ 2 Related Work and Contributions
In this section, we present a comprehensive overview of relevant literature for this work, focusing on the following central topics: Spiking Neural Networks (Section 2.1); Forward-only Learning Algorithms (Section 2.2); Out-of-Distribution Detection (Section 2.3); and Explainability and Attribution (Section 2.4). Finally, Section 2.5 details the contribution that this papers offer with respect to the current literature.
@@ 2.1 Spiking Neural Networks
SNNs constitute a special category of artificial neural networks inspired by biological models, aiming to replicate the communication patterns found in real neurons within the brain. Unlike traditional approaches, primarily driven by Analog Neural Networks (ANNs) that rely on continuous activation values to forward information, spiking models operate by using discrete pulses of information, known as spikes. The internal dynamics of the neuron, determined by the chosen neural model, dictate when spikes occur and are typically presented as a dynamical system that provides a mathematical approximation of the behavior observed in real neurons (Yamazaki et al., 2022). As such, depending on the degree of biological plausibility desired, models can be more accurate but computationally expensive (e.g., Hodgkin-Huxley), or they can satisfy multiple properties observed in real neurons while maintaining low computational complexity (e.g., Integrate-and-Fire models).
In this work, we employ the Leaky Integrate-and-Fire (LIF) neural model, which is known for preserving the standard action potential dynamics and being easily computable (Abbott, 1999; Gerstner et al., 2014). Under this model, the membrane potential $U$ of the neuron is tracked and updated in response to the external stimulus $I_{\text{in}}$. This activity decays over time by a decay factor $\tau_{t}$, creating the leaky dynamic that gives the model its name. The dynamics of the neuron can be described as follows:
$U(t)=\tau_{t}\cdot U(t-1)+R\cdot I_{\text{in}}(t),$ (1)
where $U(t-1)$ represents the membrane potential at the previous time step, and $R$ is a scaling factor of the input.
Once the membrane potential reaches a threshold $\theta_{\text{thr}}$, a spike is generated and forwarded to subsequent neurons. At this point, the membrane potential is reset to zero, resulting in the following equations:
@@ 2.2 Forward-only Learning Algorithms
Backpropagation stands as the most common algorithm for training artificial neural networks due to its high adaptability and precision on different architectures and loss objectives. However, the need for differentiable activations when creating the network poses a clear limitation in spiking models, which use non-differentiable neural models. To address this, the most commonly employed technique relies on surrogate gradients, which involve replacing the original neural model with a differentiable approximation during the backward pass, so that the gradient value remains close to the expected gradient direction (Neftci et al., 2019).
@@
Recently, the Forward-Forward Algorithm was introduced as a forward-only alternative to backpropagation, advocating for a biologically-motivated heuristic in which the backward pass is replaced by an additional contrastive forward pass (Hinton, 2022). In this second pass, the model is presented with an adversarial sample, which is used to contrastively learn to distinguish between samples from the training set (which we denote by positive or $D_{\oplus}$) and synthetically created data (resp. negative or $D_{\ominus}$). As it follows from its forward-only nature, each layer is trained independently, using a layer-specific loss that solely employs local information. To formulate these local updates, Hinton proposed the use of a goodness score, which offers a numeric measurement related to the likelihood of a latent vector belonging to the positive data distribution. This goodness function $G:\mathbb{L}\rightarrow\mathbb{R}$ was defined as the squared Euclidean norm of the latent vector, where $\mathbb{L}$ denotes the latent space of the given layer. Under this formulation, the goodness score serves as a measure of the layer’s activity, thereby expecting highly active latent vectors when presented with positive samples, and nearly dormant vectors when encountering a negative instance. By combining this goodness score with a probability function $P:\mathbb{R}\rightarrow[0,1]$, mapping the goodness domain into a probability range, the network can be trained by using the standard Binary Cross Entropy between positive and negative binary states. In the original work, this function was computed using a logistic function, modulated by a threshold value $\theta$ to shift the scores closer to zero when presented with near-zero goodness scores. This probability is formally denoted as $P(x;\theta)=\text{sigmoid}(x-\theta)$.
@@
As previously stated, each learning step involves two contrastive forward passes, a positive pass and a negative pass. In its supervised formulation, positive instances are created by embedding the corresponding label of the input as a one-hot encoded vector into the corner of the base input image. Analogously, negative data is created by taking samples from the training dataset and embedding random labels that do not align with their corresponding class. Using this positive and negative generation method, the inference process computes the goodness score of the samples with respect to all labels, selecting the label with the highest goodness score as the predicted class. To enhance this process, Lee and Song (2023) refined the label embedding process by associating each class with a random binary vector, which is then appended to each input, removing the overlap between the embedded labels and the data, achieving enhanced classification accuracy.
Despite its clear biological motivation, research on FFA is still heavily conducted in ANNs, with SNNs being explored in only a handful of papers. The first spiking adaptation was developed by Ororbia (2023), who proposed replacing the learning rule with a Hebbian or STDP update through time, modulated by a positive/negative reward signal. Following this work, Merkel and Ororbia (2024) provided experimental evidence that Ororbia’s biological adaptation yields positive results when training networks on physical neuromorphic chips. In a parallel effort, Terres-Escudero et al. (2024) studied the relationship between FFA’s formulation and neo-Hebbian frameworks, demonstrating how FFA represents a specific branch of contrastive neo-Hebbian learning rules. Nevertheless, research conducted with ANNs has highlighted ways to improve the method’s accuracy and revealed the geometric structure that emerges in trained networks (Tosato et al., 2023; Ororbia and Mali, 2023). Studies from Tosato et al. (2023) or Yang (2023) consistently emphasize the characteristic latent topology produced by this learning heuristic, characterized by sparse latent activity and high neural specialization, resulting in a landscape where class clusters are distinctly separated by clear boundaries.
@@ 2.3 Out-of-Distribution Detection
OoD Detection plays a crucial role in assessing the security of production-ready AI pipelines, as it ensures that models aren’t exposed to inputs that could produce unpredictable (and potentially dangerous) behavior. These algorithms operate by defining a scoring function that measures the likelihood of a sample from belonging to the ID dataset, which is then used as a filter to remove the examples that do not satisfy a user-defined likelihood threshold. The construction of these functions has been explored by using two different properties of the networks: their output probability distribution, and their internal latent space. Given the absence of a classification layer on FFA, this paper focuses on developing distance-based techniques, which provide with a distance function to the latent space to measure how far samples OoD lie with respect to the training distribution (Yang et al., 2021). Similarly, considering the lack of research on OoD Detection in forward-only methods, we offer a comparative perspective by choosing methods trained using backpropagation.
Among OoD Detection benchmarks, ODIN stands as the most widely-used classification-based method due to its accuracy and reduced computational costs. This algorithm operates by comparing the final classification distribution when scaling the softmax temperature and introducing small, gradient-based input perturbations, thereby making OoD samples produce lower confidence scores (Liang et al., 2017). Within distance-based methods, Bergman et al. (2020) proposed the deep k-nearest neighbor algorithm to measure the distance between input latent vectors and previously computed ID representatives, using these distances to detect anomalies. Building upon this work, Sun et al. (2022) extended the application of deep k-nearest neighbor distances for OoD detection, highlighting the importance of leveraging geometric properties in the latent space of trained networks for OoD detection tasks.
It is important to acknowledge that the OoD field has only recently been introduced into SNNs, with Martinez-Seras et al. (2023) offering the first fully spiking OoD detection algorithm. In their work, they introduce the Spike Count Pattern (SPC) algorithm, which characterizes spike patterns across layers to identify whether a given input aligns with the latent behavior observed with ID data. In order to produce a fair comparison between works, and given the absence of other spiking OoD papers, we rely on an equivalent experimental setup as them.
2.4 Interpretability
It is well acknowledged that neural networks operate as black boxes, achieving high performance across multiple tasks while offering little to no insight into the heuristics followed to obtain their outputs. The term Explainable AI was recently coined to group all techniques focused on developing methods aimed at offering humanly understandable explanations for these inner mechanisms (Arrieta et al., 2020). Among the methods in this domain, Interpretability techniques aim at providing heuristics for interpreting the inner behavior of a model. Within visual data, Grad-CAM-like methods stands as one of the most impactful techniques, using gradient information extracted of the last convolutional layer to detect areas of the input with the highest impact on the prediction of a specific class (Selvaraju et al., 2017). Nevertheless, due to its reliance on gradient information to obtain these attribution maps, such techniques remain elusive for spiking systems, where only a handful of methods have been proposed (Kim and Panda, 2021; Bitar et al., 2023).